OpenStudy (anonymous):

Derive a formula for the product of the solutions of the quadratic eqaution ax^2 + bx + c in terms of the coefficients

OpenStudy (jamesj):

Surely your teacher showed you this -- is it not in your notes or in your textbook?

OpenStudy (anonymous):

No it is not

OpenStudy (jamesj):

If not, then this geeky website has it for you: http://www.mathsisfun.com/algebra/quadratic-equation-derivation.html

OpenStudy (anonymous):

I know how to derive the equation but I don't understand how to do it for the product of its solutions

OpenStudy (jamesj):

Oh sorry, I misread your question.

OpenStudy (amistre64):

completeing the square is the long hand version of the quad formula

OpenStudy (anonymous):

haha that's okay

OpenStudy (anonymous):

or the sum of its solutions for that matter

OpenStudy (jamesj):

Now take the two solutions and multiply them together \[\frac{-b + \sqrt{b^2-4ac}}{2a} \ . \ \frac{-b - \sqrt{b^2-4ac}}{2a} = \frac{b^2 - (b^2 - 4ac)}{4a^2} = \ ....\]

OpenStudy (anonymous):

OHHHH!!!!! That makes so much more sense thank you.

OpenStudy (amistre64):

\[ax^2+bx+c=0\] \[x^2+\frac{b}{a}x=-\frac{c}{a}\] \[x^2+\frac{b}{a}x+(\frac{b}{2a})^2=-\frac{c}{a}+(\frac{b}{2a})^2\] \[\left(x+\frac{b}{2a}\right)^2=-\frac{c}{a}+\frac{b^2}{4a^2}\] \[\left(x+\frac{b}{2a}\right)^2=-\frac{4ac+b^2}{4a^2}\] \[x+\frac{b}{2a}=\pm\sqrt{-\frac{4ac+b^2}{4a^2}}\] \[x+\frac{b}{2a}=\pm\frac{\sqrt{-4ac+b^2}}{2a}\] \[x=-\frac{b}{2a}\pm\frac{\sqrt{-4ac+b^2}}{2a}\] \[x=-\frac{b}{2a}\pm\frac{\sqrt{4ac+b^2}}{2a}\] \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

no I know how to do that, I need a formula for the sum of its two solutions

OpenStudy (amistre64):

.... i dropped my negative in the next to last step

OpenStudy (jamesj):

So use the quadratic formula and add the two solutions together. You'll see it simplifies nicely.

OpenStudy (anonymous):

I got -2b over 2a, is that correct?

OpenStudy (jamesj):

or just -b/a, yes.

OpenStudy (anonymous):

yes, thanks so much, it makes a lot of sense now, what would be correct for the product of the solutions?

OpenStudy (jamesj):

Tell me what you think it is and and I'll tell you if it's right.

OpenStudy (anonymous):

-4C/4A?

OpenStudy (jamesj):

Not quite ...

OpenStudy (anonymous):

should "b" be in there?

OpenStudy (jamesj):

no

OpenStudy (anonymous):

-4ac/4a^2?

OpenStudy (anonymous):

NO!, not negative on top

OpenStudy (anonymous):

c over a!

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

YES!

OpenStudy (jamesj):

Now I'm going to show you another way.

OpenStudy (anonymous):

wow thanks man your awesome

OpenStudy (jamesj):

Suppose ax^2 + bx + c = 0 is a quadratic equation (and hence a is not zero) Then x^2 + (b/a) x + (c/a) = 0

OpenStudy (jamesj):

Let p and q be the roots of this equation. Then (x - p)(x - q) = 0

OpenStudy (jamesj):

i.e., x^2 - (p+q) x + pq = 0

OpenStudy (anonymous):

yes, I follow

OpenStudy (jamesj):

Comparing the two forms of the equation now, we see that p + q = -b/a and pq = c/a

OpenStudy (jamesj):

The same results we just derived.

OpenStudy (anonymous):

oh my gosh, the exact two we just found! That is so cool!

OpenStudy (jamesj):

It is cool. Maths works.

OpenStudy (anonymous):

haha right, thanks again, I have to go now, but you helped me soooo much!

OpenStudy (jamesj):

sure thing. good luck.